In a moving boundary experiment 0.01M Hcl solution was treated on a lithium chloride solution. The tube used had a diameter of 1cm .when a current of 11.0milliametre was passed for 20minutes ,the H+ ions ,Li+ ions boundary moved through 13.9cm . calculate the transport number of H+ and Cl- ions in the Hcl solution

1
Expert's answer
2022-03-30T16:08:03-0400

By Faraday's laws of electrolysis:

m = k * I * t;

M(HCl) = 36.5 g/mol

F = 96485.3 V/mol'

k(HCl) = M/F = 36.5/96485.3 = 0.00038 g/C;

I = 11.0 mA = 0.011 A;

t = 20 m = 1200 s;

m(HCl) = k * I * t = 0.00038 * 0.011 * 1200 = 0.005 g

l(length of migration) = 13.9 sm;

d(diameter of tube)

A(cross-section of tube) = "\pi" * R2 = "\pi" * (d/2)2 = 3.14 * (1/2)2 = 0.785 sm2;

C(HCl) = 0.01 M;

Q(charge passed, deposited H+) = m/M = 0.005/1 = 0.005;

tH+ = l * A * C/(1000 * Q) = 13.9 * 0.785 * 0.01/(1000 * 0.005) = 0.0218;

tH+ + tCl- = 1

tCl- = 1 - tH+ = 1 - 0.0218 = 0.9782.